# NarrativeBridge: Enhancing Video Captioning with Causal-Temporal Narrative

### v2

@@ 5 Limitations
While our work represents a significant step forward in causal-temporal narrative video captioning, our method assumes a certain level of causal and temporal coherence within the video content. Highly complex or convoluted causal relationships may challenge our approach like existing approaches, necessitating further architectural enhancements for improved generalization. Also, the current work primarily focuses on understanding and generating captions that capture causal narratives and temporal dynamics. Explicit integration of spatial reasoning, multi-agent interactions, and long-term dependencies could further enhance the robustness and applicability of our approach. Despite these limitations, our work opens up new avenues for innovative applications and research directions, further solidifying the importance of causal-temporal narrative understanding in video analysis tasks.
@@ 6 Conclusion
In this paper, we introduced for the first time a Causal-Temporal Narrative (CTN) captions benchmark and proposed a novel Cause-Effect Network (CEN) tailored for causal-temporal narrative video captioning. This work finds vast applications in automated video summarization, question-answering, assistive technologies, surveillance, and educational content creation. The CTN captions benchmark provides a comprehensive testbed for evaluating video understanding models’ ability to grasp complex temporal and causal dynamics, which will be released for research purpose on https://narrativebridge.github.io/. And CEN explicitly models cause-effect relationships and temporal dynamics to generate rich, contextually relevant descriptions capturing nuanced causal-temporal narrative in videos, demonstrating significant performance improvement over SOTA methods (Section 4). NarrativeBridge lays the foundation for a paradigm shift, where models comprehend underlying causal-temporal narrative driving events, unlocking new frontiers in contextually aware human-machine interactions with video.
@@ A.2 Prompt Design Process
We then progressively improve the prompt quality, moving to a more concise set of instructions in Prompt 3 and a relatively well-structured prompt in Prompt  4. While the captions generated by these improved prompts (Figures 7 and 8) show some enhancements, they still fall short in terms of capturing the causal-temporal narrative accurately and consistently. Additionally, the LLM continues to generate multiple captions and deviate from the desired output format.
@@
These experiments highlight the significance of effective prompt design and the advantages of few-shot learning in generating high-quality CTN captions. By providing carefully crafted example captions in the prompt, we enable the LLM to better understand the desired format, style, and level of detail for the generated captions. The few-shot based prompt allows the LLM to learn from and mimic the provided examples, resulting in captions that accurately capture the causal-temporal narrative of the video content.
Crucially, the few-shot based prompt also guides the LLM to generate a single coherent caption in the specified JSON format, ensuring consistency and usability of the generated captions. In contrast, the zero-shot prompts, even when well-structured, rely solely on the LLM’s pre-existing knowledge and understanding of the task description. While the LLM can generate relevant captions to some extent, the lack of explicit examples limits its ability to produce captions with the same level of coherence, accuracy, and contextual understanding as those generated using the few-shot based prompt. Moreover, the zero-shot prompts fail to guide the LLM toward generating a single caption in the desired JSON format, resulting in multiple captions and inconsistent output structures.
@@
These findings underscore the importance of our few-shot based prompt design in the CTN caption generation process and demonstrate its superiority over zero-shot prompting approaches. By leveraging the power of few-shot learning, we enable the LLM to generate high-quality CTN captions that effectively capture the cause-effect relationships and temporal dynamics in video content, as evidenced by the results presented in Figure 1. Additionally, our few-shot based prompt ensures that the generated captions adhere to the specified JSON format, facilitating seamless integration and usability in downstream applications.
@@ A.5.4 Analysis
These findings demonstrate that our automatic generation and evaluation process, as described in Section 3.1, produces high-quality CTN captions that accurately represent the causal and temporal dynamics in video content. This human evaluation complements our automatic evaluation metrics, providing a comprehensive validation of our CTN caption benchmark.
@@ A.6 Encoder Comparison for CEN
To explore the versatility of our CEN architecture, we conduct experiments using BLIP Li et al. (2022b) encoders in place of CLIP encoders. This comparison allows us to assess the adaptability of our method to different visual encoding architectures. Table 4 presents the results of this comparison on both MSR-VTT and MSVD datasets.
@@
MSR-VTT CLIP (Ours) 27.90 49.87 15.76
@@ A.7 Comparison with NExT-QA
To highlight the differences between our CTN benchmark and the NExT-QA dataset, we provide a detailed comparison table and illustrative examples:
@@ A.8 Application of CTN Caption Generation for Labeling Unlabeled Videos
A.9 Additional Qualitative Results for CEN Architecture

### v3

@@ 5 Limitations
While our work represents a significant step forward in causal-temporal narrative video captioning, complex onvoluted causal relationships are still a challenge. This necessitates further architectural enhancements for improved generalization. Explicit integration of spatial reasoning, multi-agent interactions, and long-term dependencies could further enhance the robustness and applicability of our approach. Despite these limitations, our work opens up new avenues for innovative applications and research directions, further solidifying the importance of causal-temporal narrative understanding in video analysis tasks.
@@ 6 Conclusion
In this paper, we introduced for the first time a Causal-Temporal Narrative (CTN) captions benchmark dataset and proposed a novel Cause-Effect Network (CEN) tailored for causal-temporal narrative video captioning. This work finds vast applications in automated video summarization, question-answering, assistive technologies, surveillance, and educational content creation. The CTN captions benchmark dataset provides a comprehensive testbed for evaluating video understanding models’ ability to grasp complex temporal and causal dynamics, which will be released for research purpose on https://narrativebridge.github.io/. And CEN explicitly models cause-effect relationships and temporal dynamics to generate rich, contextually relevant descriptions capturing nuanced causal-temporal narrative in videos, demonstrating significant performance improvement over SOTA methods (Section 4). NarrativeBridge lays the foundation for a paradigm shift, where models comprehend underlying causal-temporal narrative driving events, unlocking new frontiers in contextually aware human-machine interactions with video.
@@ Acknowledgement
Ethics Statement
Our work on NarrativeBridge focuses on improving video captioning through causal-temporal narrative understanding. We use publicly available datasets (MSR-VTT and MSVD) and have cited their creators appropriately. We will release our new CTN captions benchmark dataset on our project webpage https://narrativebridge.github.io/, where we also mention the licenses of all assets used. Our research does not involve human subjects or crowdsourcing, and we do not use or curate data containing personally identifiable information or offensive content. We have read and adhered to the ethics review guidelines for ICLR submissions. We acknowledge potential ethical concerns (e.g., privacy in surveillance, risk of misleading content) but are committed to responsible development.
Reproducibility Statement
We will release the CTN captions benchmark dataset and CEN model weights upon paper publication on our project website https://narrativebridge.github.io/. We provide comprehensive information on datasets (Section 4.3.1), evaluation metrics (Section 4.1), and implementation details (Section 4.2 and Appendix A.1). We specify data splits, hyperparameters, and other key details necessary for reproducing our results. We include information about the type of GPUs and compute resources used in Section 4.2 and Appendix A.1. We properly cite all existing datasets (MSR-VTT, MSVD) and models (CLIP-ViT, Uni-VL) used in our work.
@@ A.2 Prompt Design Process
We then progressively improve the prompt quality, moving to a more concise set of instructions in Prompt 3 and a relatively well-structured prompt in Prompt  4. While the captions generated by these improved prompts (Figures 9 and 10) show some enhancements, they still fall short in terms of capturing the causal-temporal narrative accurately and consistently. Additionally, the LLM continues to generate multiple captions and deviate from the desired output format.
@@
Prompt 1 Ablations: Here, we present systematic ablation experiments analyzing the impact of individual requirements in the design of Prompt 1.
[Same as Prompt 1 but without rule 1 about using only information from the captions]
Prompt 6: Removing the grounding requirement from Prompt 1
{ "Cause": "a reckless driver speeds through a field in an unstable off-road vehicle", "Effect": "the vehicle crashes dramatically and a group celebrates with beer pong nearby" }
Figure 12: Caption generated without grounding requirement
Removing the grounding requirement results in the introduction of speculative details that were not present in the source captions. This demonstrates how the grounding requirement maintains factual accuracy.
[Same as Prompt 1 but without rule 5 about temporal consistency]
Prompt 7: Removing temporal consistency requirement from Prompt 1
{ "Cause": "people celebrate enthusiastically at a gathering", "Effect": "a car flips over while driving through an open field" }
Figure 13: Caption generated without temporal consistency requirement
The absence of the temporal consistency requirement leads to incorrect event ordering, with the celebration incorrectly positioned before the car crash. This highlights the importance of maintaining proper temporal sequence in causal narratives.
[Same as Prompt 1 but without rule 7 about word limit]
Prompt 8: Removing word limit requirement from Prompt 1
{ "Cause": "an adventurous individual decided to drive their vehicle recklessly through an open field without proper consideration for safety or stability, leading to a dangerous situation", "Effect": "the vehicle inevitably lost control and dramatically flipped over, resulting in a serious accident while nearby, a group of people were engaging in a casual game of beer pong, seemingly unaware of the incident" }
Figure 14: Caption generated without word limit requirement
Removing the word limit results in verbose descriptions with unnecessary elaboration and speculation about circumstances and motivations.
[Same as Prompt 1 but without rule 6 about plain English]
Prompt 9: Removing plain English requirement from Prompt 1
{ "Cause": "the automobile’s operator executed an ill-advised traversal of the pastoral terrain", "Effect": "culminating in vehicular inversion whilst revelry ensued in proximity" }
Figure 15: Caption generated without plain English requirement
Without the plain English requirement, the model employs unnecessarily complex vocabulary and formal language that reduces clarity and accessibility.
[Same as Prompt 1 but without rule 4 about avoiding conclusions]
Prompt 10: Removing the requirement to avoid conclusions from Prompt 1
{ "Cause": "driver’s obvious inexperience and poor judgment led to dangerous driving in field", "Effect": "predictable crash occurred while others carelessly continued their party, showing clear negligence" }
Figure 16: Caption generated without avoiding conclusions requirement
Removing the requirement to avoid conclusions leads to subjective interpretations about driver inexperience and partygoer negligence that cannot be definitively determined from the video content.
[Same as Prompt 1 but without rule 2 about relevance]
Prompt 11: Removing relevance requirement from Prompt 1
{ "Cause": "a vehicle drives through a field on a sunny day and crashes", "Effect": "people enjoy various outdoor activities including beer pong and socializing" }
Figure 17: Caption generated without relevance requirement
The absence of the relevance requirement results in the inclusion of peripheral details.
@@
This ablation analysis validates the prompt design choices and underscores the importance of maintaining all requirements for generating effective causal-temporal narrative captions. In conclusion, these findings underscore the importance of our few-shot based prompt design in the CTN caption generation process and demonstrate its superiority over zero-shot prompting approaches. By leveraging the power of few-shot learning, we enable the LLM to generate high-quality CTN captions that effectively capture the cause-effect relationships and temporal dynamics in video content, as evidenced by the results presented in Figure 1. Additionally, our few-shot based prompt along with requirements ensures that the generated captions adhere to the specified JSON format, facilitating seamless integration and usability in downstream applications.
@@ A.5.4 Analysis
These findings demonstrate that our automatic generation and evaluation process, as described in Section 3.1, produces high-quality CTN captions that accurately represent the causal and temporal dynamics in video content. This human evaluation complements our automatic evaluation metrics, providing a comprehensive validation of our CTN caption benchmark dataset.
@@ A.6 Encoder Comparison for CEN
To explore the versatility of our CEN architecture, we conduct experiments using BLIP Li et al. (2022b) encoders in place of CLIP encoders. This comparison allows us to assess the adaptability of our method to different visual encoding architectures. Table 3 presents the results of this comparison on both MSVD-CTN and MSRVTT-CTN datasets.
@@
MSVD-CTN CLIP (Ours) 31.46 63.51 19.25
BLIP 28.79 57.23 18.06
MSRVTT-CTN CLIP (Ours) 27.90 49.87 15.76
@@ A.7 Comparison with NExT-QA
To highlight the differences between our CTN benchmark dataset and the NExT-QA dataset, we provide a detailed comparison table and illustrative examples:
@@ A.8 Application of CTN Caption Generation for Labeling Unlabeled Videos
A.9 LLM-Based Evaluation of generated CTN captions
We conduct additional evaluations on MSVD-CTN and MSRVTT-CTN (combined) to evaluate further the quality of generated CTN captions, using Llama-3.2-3B-Instruct in the default configuration. Our evaluation focused on two key aspects:
1. Temporal Order Analysis
  • Input: Ground truth and generated captions
  • Task: LLM compares the temporal sequence of events
  • Score: Binary (1: correct sequence, 0: incorrect)
2. Causal Chain Analysis
  • Input: Ground truth and generated captions
  • Task: LLM evaluates preservation of cause-effect relationships
  • Score: Binary (1: preserved, 0: not preserved)
Table 5 shows the performance comparison between CEN and the best baseline (GIT).
Table 5: LLM-based evaluation results comparing CEN with GIT
Model Temporal Order (%) Causal Chain (%)
CEN (Ours) 81.2 84.5
GIT 52.1 48.3
We further analyze all ablations of Table 6 using the same LLM-based metrics, as shown in Table 6.
Table 6: LLM-based evaluation results for ablation study
Method Temporal Order (%) Causal Chain (%)
CEN (Full) 81.2 84.5
Ecombined 72.4 75.6
w/o FT - single CLIP 58.3 54.2
w/o FT - Two CLIP 63.7 61.8
Only Ecause 73.9 76.5
Only Eeffect 74.8 77.2
Zero Shot X 55.2 51.7
Fine-tune X 79.4 82.1
Following are the key findings:
  • CEN outperforms GIT by 29.1% in temporal and 36.2% in causal accuracy
  • Dual encoder architecture provides 15% improvement over single encoder
  • Fine-tuning improves temporal understanding by  20% compared to frozen models
A.10 Additional Qualitative Results for CEN Architecture